OpenStudy (mony01):

Calculate the following:

OpenStudy (mony01):

\[\int\limits x^3\exp(5x^4)dx\]

OpenStudy (kirbykirby):

Let \(u = 5x^4\) \(du = 5(4)x^3\, dx\implies \dfrac{du}{20}=x^3\, dx\)

OpenStudy (kirbykirby):

The integral becomes very simple then \[ \frac{1}{20}\int e^u\, du\]

OpenStudy (mony01):

thanks!!